OpenStudy (anonymous):

The equilibrium constant for the reaction of carbon monoxide with water is 1.845. If 1.00 mol of each reactant is placed in a 2.00 L container, what is the concentration of carbon dioxide at equilibrium? CO(g) + H2O(g) CO2(g) + H2(g)

OpenStudy (cuanchi):

0.287983 M = 0.288 M CO + H20 ----> CO2 + H2 I 0.5M 0.5M 0.0 0.0 C -x -x +x +x E 0.5-x 0.5-x x x \[Keq = \frac{ [CO _{2}] [H2] }{ [CO] [H _{2} O]} = \frac{ x ^{2} }{ (0.5-x)^{2} }= 1.845\] calculate the value for x \[x = \frac{ 0.5 \sqrt{1.845} }{ 1+ \sqrt{1.845} } = 0.2879= 0.288 M\]